OpenStudy (anonymous):

The government injects $90 billion into the economy. Each recipient spends 90% of the dollars received. In turn, the secondary recipients spend 90% of the dollars they receive, etc. Use the expansion of 1/(1-x) in powers of x to come up with a clean measurement of the total spending that results from the $90 billion injection.

OpenStudy (anonymous):

\[90\times .9+90\times .9^2+90\times .9^3+...\] \[10(1+.9+.9^2+.9^3+...\] \[\frac{10}{1-.9}=\frac{10}{.1}=100\]

OpenStudy (anonymous):

ahh makes sense! thank you so much!

OpenStudy (anonymous):

yw just hope it is right